OpenStudy (anonymous):

can someone help me factor x^2+4x-16?

OpenStudy (anonymous):

it's really not factorable

OpenStudy (anonymous):

\[x = \frac{-4 \pm \sqrt{16 + 4.16}}{2}\] \[x = -2 + 2\sqrt{5},-2 - 2\sqrt5\]

OpenStudy (anonymous):

unless you use the quadratic formula

OpenStudy (anonymous):

ah yeah it is but through myininaya method

OpenStudy (anonymous):

do you want to see that

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

okay then ax^2 + bx + c =0 (b/2 + z)(b/2 - z) = a*c b^2/4 -z^2 = ac z^2 = b^2/4 - ac z = sqrt( b^2 - 4ac)/2 this gives z = 4sqrt5/2 = 2sqrt5 then put b/2+z and b/2 -z inplace of bx

OpenStudy (anonymous):

wait a minute i will post a link

OpenStudy (amistre64):

x^2+4x-16 = 0 ; -16 x^2 +4x + ___ =16 + ___ ; complete the square x^2 +4x +4 = 16+4 ; compact it (x+2)^2 = 20 ; now we can either go further or simply see that sqrt(20) is what we really need so subtract out the 2 -2 + sqrt(20) -2 - sqrt(20) ; now put into factored form .................................................... (x +2 - sqrt(20)) (x -2 -sqrt(20))

OpenStudy (amistre64):

mighta typoed it .... dbl chk the format :)